OpenStudy (anonymous):

What is the range of the function f(x) = 2x-5?

OpenStudy (hockeychick23):

the range is all your y values once you plug in values for the equation for example maybe you choose to use 0,1,2,3 as your x values. Now you have to plug each one of these values in. f(x)=2(0)-5 f(x)= -5 f(x)=2(1)-5 f(x)=-3 f(x)= 2(2)-5 f(x)=-1 f(x)= 2(3)-5 f(x)=1 so -5,-3,-1 are your y values when x is 0,1,2, and 3 your y values are your range and your x values are your domain